An architect designs a rectangular flower garden such that the width is exactly two-thirds of the length. If 260 feet of antique picket fencing are to be used to enclose the garden, find the dimension of the garden
Answer by jorel1380(3719)

W=2/3L
2L+2W=260
2L+2(2/3L)=260
6/3L+4/3L=260
10/3L=260
10L=780
L=78ft.
W=52ft..